What change in velocity would produce a body of mass 4Kg if a constant force of 16N acts on it for 2 seconds
1
Expert's answer
2021-10-04T15:30:42-0400

Applying the second Newton's law (for constant force), obtain:


"m\dfrac{\Delta v}{\Delta t} = F"

where "m = 4kg, \Delta t =2s, F=16N" and "\Delta v" is the required change in velocity.

Thus, obtain:


"\Delta v = \dfrac{F\Delta t}{m}\\
\Delta v = \dfrac{16N\cdot 2s}{4kg} = 8m/s"

Answer. 8m/s.